# Searching for Pulsars, Magnetars, and Fast Radio Bursts in the Sculptor Galaxy using MeerKAT

### v1

@@ Abstract
An increasing number of fast radio bursts (FRBs) have been detected in the last few decades. FRBs are classified by their one-off or multiple short-duration bursts (Lorimer et al., 2007; Thornton et al., 2013; Spitler et al., 2016), and it is accepted that FRBs are extragalactic events due to the high associated dispersion measure (DM) that put them well beyond the MW limits (e.g., Platts et al. 2019). The mechanisms responsible for repeating and non-repeating bursts are still a matter of debate. With the detection and localisation of more bursts, constraints can be placed on their emission mechanism (e.g., CHIME/FRB Collaboration et al. 2021). A possible explanation for non-repeating FRBs is that they are an after-effect of cataclysmic events (e.g., Marnoch et al. 2020). On the contrary, repeating FRBs should have a non-cataclysmic origin (Spitler et al., 2016). A recent detection of FRB-like bursts was made from the Soft-Gamma Repeater (SGR) J1935+2154, which is a Galactic magnetar. This suggests a strong connection between magnetar emission and at least some repeating FRBs (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Camilo et al., 2006a). Furthermore, SGRs are associated with regions of recent star-forming and supernovae (Gaensler, 2004). Known FRB host galaxies exhibit a wide range of compositions, with masses in the range – and star-formation rates (SFRs) ranging from – (Bhandari et al., 2022). Mannings et al. (2021) found that all FRBs located in hosts with spiral structures occur near or on a spiral arm.
@@
The Nearby Galaxies programme within the TRansients And PUlsars with MeerKAT (TRAPUM222https://www.trapum.org) Large Survey Project (Stappers & Kramer, 2016), has as one of its aims the discovery of new pulsars and fast transients outside of the MW galaxy, and to investigate their properties in relation to those of their host galaxy. One of the targets of this survey is NGC 253, also known as the Sculptor Galaxy. It is located in the southern sky far below the Galactic Plane, at a Galactic latitude of and a distance of Mpc (Rekola et al., 2005). It is a promising target for finding pulsars, due to its population of massive stars and its similarity to the MW. It is classified as a barred spiral galaxy and is one of the nearest galaxies undergoing a nuclear starburst. It is one of the brightest starburst galaxies in the Southern Hemisphere, with detections of numerous supernova remnants (SNRs) within the starburst region, including one optically identified supernova (SNR 1940E, Chhetri et al. 2018; Zwicky 1941) close to the central region of the galaxy. The unobscured stellar population is consistent with ages Myr (Kornei & McCrady, 2009; Davidge, 2016). The centre hosts a population of more than a dozen super star clusters (SSC) that are still in the formative stages. A near-infrared photometry study of this galaxy revealed 181 star clusters in the central 600 pc (Camperi et al., 2022). Nine of these clusters have masses in the range of 10 M⊙ and ages less than 7 Myr, with the remaining clusters being older and less massive. This is interesting since these clusters can provide insights into the star formation history of this galaxy. NGC 253 has a total mass of (Karachentsev et al., 2021), which is about a factor of two less than the total mass of the MW. Forbes et al. (1993) suggested that the central 6 of NGC 253 contains about 24,000 O-type stars and that SNRs are located throughout the nuclear region. The nucleus has an SFR of approximately , which accounts for about 70 per cent of the rate of the entire galaxy (Wik et al., 2014). The central SFR of NGC 253 is 30–40 times higher than that of the central region of the MW (Longmore et al., 2013; Barnes et al., 2017; Mills et al., 2021). A high SFR is consistent with a young stellar population (Duarte Puertas et al., 2022), which in turn is likely to contribute to a large SNR population. Furthermore, Wik et al. (2014) found that X-ray emission above 10 keV were concentrated in the central 100 of NGC 253, produced by three nuclear sources, an off-nuclear ultraluminous X-ray source and a pulsar candidate, which could be an extremely luminous X-ray pulsar. NGC 253 is furthermore the first galaxy to be associated with two magnetar giant flares (MGFs), (Gamma-Ray Burst) GRB 180128A and GRB 200415A (Trigg et al. 2023; Svinkin et al. 2021). GRB 200415A was observed on April 15, 2020, located in an area of 20 arcmin2 that overlaps the central region of NGC 253 (Svinkin et al., 2021). GRB 180128A occured 808 days before GRB 200415A on January 28, 2018, and was found in archival data after the discovery of GRB 200415A. However, it is still unclear whether these bursts originated from the same magnetar (Trigg et al., 2023). The formation channel of magnetars is consistent with core-collapse supernovae (Heintz et al., 2020; Bochenek et al., 2021). Thus, the high SFR, the presence of SNRs, the high energy X-ray sources, young star clusters, and associated MGFs support the likelihood that NGC 253 contains a young NS/pulsar population. It is also interesting to note that young pulsars sometimes emit giant radio pulses (Lundgren et al., 1995; Geyer et al., 2021), which would enable us to discover distant radio pulsars that might otherwise not be detectable (McLaughlin & Cordes, 2003). The probable presence of at least one magnetar in this galaxy also means that there may be sources responsible for emitting FRB-like pulses, like those seen from SGR J1935-2154.
@@ 3 Search sensitivity and upper limits
where S/N is the spectral S/N threshold of 9.5, is the correction factor due to digitization. Since the loss of telescope sensitivity due to 8-bit digitization is minimal, we have = 1 (Kouwenhoven & Voûte 2001).A conversion efficiency factor for converting from spectral S/N to folded S/N (Morello et al., 2018). The system temperature is the sum of the receiver temperature ( = 18 K, Bailes et al. 2020) and the sky temperature contribution 4.24 K (Zheng et al., 2017)111111http://github.com/jeffzhen/gsm2016 in the direction of NGC 253. The telescope gain is G = 1.925 K Jy-1 for the core array in coherent mode (G = 0.34 for incoherent mode with 61 antennas). Lastly, the number of polarisations is , the bandwidth MHz, and the integration time = 14085.30 s. We assume a pulsar duty cycle of , which is the median intrinsic pulsar duty cycle excluding millisecond pulsars and pulsars in globular clusters, rotating radio transients, magnetars and binary pulsars, taken from the ATNF catalogue (Manchester et al., 2005), and a spectral index of (Jankowski et al., 2018). Thus, the minimum detectable flux density for our periodicity search is = 4.4 Jy at the centre of the incoherent beam and of a coherent beam for a pulsar with a period of 100 ms and above at a DM = 250 pc cm-3. At lower spin periods, factors such as dispersion smearing, sampling time, and scattering influence the detectability of such short period pulsars. The corresponding pseudo-luminosity limit, which is calculated using
@@ 4 Discussion & Conclusion
Moving to single-pulse searches, as seen in Figure 2, we were sensitive to FRB and FRB-like pulse detections. We derived a single-pulse flux density limit of approximately 62 mJy for our FRB search in the direction of NGC 253. This limit is orders of magnitude lower than the flux density of previously detected FRBs. Since two MGFs are associated with NGC 253 (Trigg et al., 2023; Svinkin et al., 2021), we would expect to be able to observe bursts similar to those of SGR J1935+2154 in the MW, should they have occurred. However, our non-detection can potentially be due to the epoch of observation, since these events are unpredictable and sporadic. It can also be due to unfavourable orientations. Transient radio emission, lasting for years have also been seen from magnetars, but none has been associated with MGFs, since these events are rare and not much radio follow up has been possible (Camilo et al., 2006b; Levin et al., 2019)
@@ Acknowledgements
We thank the anonymous referee for their constructive feedback that helped improve the quality of the paper. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation (NRF), an agency of the Department of Science and Innovation (DSI). SARAO acknowledges the ongoing advice and calibration of GPS systems by the National Metrology Institute of South Africa (NMISA) and the time space reference systems department of the Paris Observatory. TRAPUM observations used the FBFUSE and APSUSE computing clusters for data acquisition, storage and analysis. These clusters were designed, funded and installed by the Max-Planck-Institut für Radioastronomie and the Max-Planck Gesellschaft. This work is based on research supported in part by the NRF. C.V. acknowledges that opinions, findings and conclusions or recommendations expressed in any publication generated by the NRF-supported research are those of the author(s), and that the NRF accepts no liability whatsoever in this regard. H.H. acknowledges research support in part by the NRF of South Africa (Grant number 142 859), and support through the National Astrophysics and Space Science Program (NASSP), funded by DSI-NRF. R.P.B. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 715051; Spiders) This research has made use of hips2fits,121212https://alasky.cds.unistra.fr/hips-image-services/hips2fits a service provided by CDS, the SIMBAD database, operated at CDS, Strasbourg, France (Wenger et al., 2000), NASA’s Astrophysics Data System Bibliographic Services, and the ATNF pulsar catalogue version 2.3.0. Figure 1 was made using the APLpy (Robitaille & Bressert, 2012; Robitaille, 2019), which is an open-source astronomy plotting Python package.

### v2

@@ Abstract
An increasing number of fast radio bursts (FRBs) have been detected in the last few decades. FRBs are classified by their one-off or multiple short-duration bursts (Lorimer et al., 2007; Thornton et al., 2013; Spitler et al., 2016), and it is accepted that FRBs are extragalactic events due to the high associated dispersion measure (DM) that put them well beyond the MW limits (e.g., Platts et al. 2019). The mechanisms responsible for repeating and non-repeating bursts are still a matter of debate. With the detection and localisation of more bursts, constraints can be placed on their emission mechanism (e.g., CHIME/FRB Collaboration et al. 2021). A possible explanation for non-repeating FRBs is that they are an after-effect of cataclysmic events (e.g., Marnoch et al. 2020). On the contrary, repeating FRBs should have a non-cataclysmic origin (Spitler et al., 2016). A recent detection of FRB-like bursts was made from the Soft-Gamma Repeater (SGR) J1935+2154, which is a Galactic magnetar. This suggests a strong connection between magnetar emission and at least some repeating FRBs (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020; Camilo et al., 2006). Furthermore, SGRs are associated with regions of recent star-forming and supernovae (Gaensler, 2004). Known FRB host galaxies exhibit a wide range of compositions, with masses in the range – and star-formation rates (SFRs) ranging from – (Bhandari et al., 2022). Mannings et al. (2021) found that all FRBs located in hosts with spiral structures occur near or on a spiral arm.
@@
The Nearby Galaxies programme within the TRansients And PUlsars with MeerKAT (TRAPUM222https://www.trapum.org) Large Survey Project (Stappers & Kramer, 2016), has as one of its aims the discovery of new pulsars and fast transients outside of the MW galaxy, and to investigate their properties in relation to those of their host galaxy. One of the targets of this survey is NGC 253, also known as the Sculptor Galaxy. It is located in the southern sky far below the Galactic Plane, at a Galactic latitude of and a distance of Mpc (Rekola et al., 2005). It is a promising target for finding pulsars, due to its population of massive stars and its similarity to the MW. It is classified as a barred spiral galaxy and is one of the nearest galaxies undergoing a nuclear starburst. It is one of the brightest starburst galaxies in the Southern Hemisphere, with detections of numerous supernova remnants (SNRs) within the starburst region, including one optically identified supernova (SNR 1940E, Chhetri et al. 2018; Zwicky 1941) close to the central region of the galaxy. The unobscured stellar population is consistent with ages Myr (Kornei & McCrady, 2009; Davidge, 2016). The centre hosts a population of more than a dozen super star clusters (SSC) that are still in the formative stages. A near-infrared photometry study of this galaxy revealed 181 star clusters in the central 600 pc (Camperi et al., 2022). Nine of these clusters have masses in the range of 10 M⊙ and ages less than 7 Myr, with the remaining clusters being older and less massive. This is interesting since these clusters can provide insights into the star formation history of this galaxy. NGC 253 has a total mass of (Karachentsev et al., 2021), which is about a factor of two less than the total mass of the MW. Forbes et al. (1993) suggested that the central 6 of NGC 253 contains about 24,000 O-type stars and that SNRs are located throughout the nuclear region. The nucleus has an SFR of approximately , which accounts for about 70 per cent of the rate of the entire galaxy (Wik et al., 2014). The central SFR of NGC 253 is 30–40 times higher than that of the central region of the MW (Longmore et al., 2013; Barnes et al., 2017; Mills et al., 2021). A high SFR is consistent with a young stellar population (Duarte Puertas et al., 2022), which in turn is likely to contribute to a large SNR population. Furthermore, Wik et al. (2014) found that X-ray emission above 10 keV were concentrated in the central 100 of NGC 253, produced by three nuclear sources, an off-nuclear ultraluminous X-ray source and a pulsar candidate, which could be an extremely luminous X-ray pulsar. NGC 253 is furthermore the first galaxy to be associated with two magnetar giant flares (MGFs), (Gamma-Ray Burst) GRB 180128A and GRB 200415A (Trigg et al. 2024; Svinkin et al. 2021). GRB 200415A was observed on April 15, 2020, located in an area of 20 arcmin2 that overlaps the central region of NGC 253 (Svinkin et al., 2021). GRB 180128A occured 808 days before GRB 200415A on January 28, 2018, and was found in archival data after the discovery of GRB 200415A. However, it is still unclear whether these bursts originated from the same magnetar (Trigg et al., 2024). The formation channel of magnetars is consistent with core-collapse supernovae (Heintz et al., 2020; Bochenek et al., 2021). Thus, the high SFR, the presence of SNRs, the high energy X-ray sources, young star clusters, and associated MGFs support the likelihood that NGC 253 contains a young NS/pulsar population. It is also interesting to note that young pulsars sometimes emit giant radio pulses (Lundgren et al., 1995; Geyer et al., 2021), which would enable us to discover distant radio pulsars that might otherwise not be detectable (McLaughlin & Cordes, 2003). The probable presence of at least one magnetar in this galaxy also means that there may be sources responsible for emitting FRB-like pulses, like those seen from SGR J1935-2154.
@@ 3 Search sensitivity and upper limits
where S/N is the spectral S/N threshold of 9.5, is the correction factor due to digitization. Since the loss of telescope sensitivity due to 8-bit digitization is minimal, we have = 1 (Kouwenhoven & Voûte 2001).A conversion efficiency factor for converting from spectral S/N to folded S/N (Morello et al., 2020). The system temperature is the sum of the receiver temperature ( = 18 K, Bailes et al. 2020) and the sky temperature contribution 4.24 K (Zheng et al., 2017)111111http://github.com/jeffzhen/gsm2016 in the direction of NGC 253. The telescope gain is G = 1.925 K Jy-1 for the core array in coherent mode (G = 0.34 for incoherent mode with 61 antennas). Lastly, the number of polarisations is , the bandwidth MHz, and the integration time = 14085.30 s. We assume a pulsar duty cycle of , which is the median intrinsic pulsar duty cycle excluding millisecond pulsars and pulsars in globular clusters, rotating radio transients, magnetars and binary pulsars, taken from the ATNF catalogue (Manchester et al., 2005), and a spectral index of (Jankowski et al., 2018). Thus, the minimum detectable flux density for our periodicity search is = 4.4 Jy at the centre of the incoherent beam and of a coherent beam for a pulsar with a period of 100 ms and above at a DM = 250 pc cm-3. At lower spin periods, factors such as dispersion smearing, sampling time, and scattering influence the detectability of such short period pulsars. The corresponding pseudo-luminosity limit, which is calculated using
@@ 4 Discussion & Conclusion
Moving to single-pulse searches, as seen in Figure 2, we were sensitive to FRB and FRB-like pulse detections. We derived a single-pulse flux density limit of approximately 62 mJy for our FRB search in the direction of NGC 253. This limit is orders of magnitude lower than the flux density of previously detected FRBs. Since two MGFs are associated with NGC 253 (Trigg et al., 2024; Svinkin et al., 2021), we would expect to be able to observe bursts similar to those of SGR J1935+2154 in the MW, should they have occurred. However, our non-detection can potentially be due to the epoch of observation, since these events are unpredictable and sporadic. It can also be due to unfavourable orientations. Transient radio emission, lasting for years have also been seen from magnetars, but none has been associated with MGFs, since these events are rare and not much radio follow up has been possible (Camilo et al., 2006; Levin et al., 2019)
@@ Acknowledgements
We thank the anonymous referee for their constructive feedback that helped improve the quality of the paper. The MeerKAT telescope is operated by the South African Radio Astronomy Observatory, which is a facility of the National Research Foundation (NRF), an agency of the Department of Science and Innovation (DSI). SARAO acknowledges the ongoing advice and calibration of GPS systems by the National Metrology Institute of South Africa (NMISA) and the time space reference systems department of the Paris Observatory. TRAPUM observations used the FBFUSE and APSUSE computing clusters for data acquisition, storage and analysis. These clusters were designed, funded and installed by the Max-Planck-Institut für Radioastronomie and the Max-Planck Gesellschaft. This work is based on research supported in part by the NRF. C.V. acknowledges that opinions, findings and conclusions or recommendations expressed in any publication generated by the NRF-supported research are those of the author(s), and that the NRF accepts no liability whatsoever in this regard. H.H. acknowledges research support in part by the NRF of South Africa (Grant number 142 859), and support through the National Astrophysics and Space Science Program (NASSP), funded by NRF. R.P.B. acknowledges support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 715051; Spiders) This research has made use of hips2fits,121212https://alasky.cds.unistra.fr/hips-image-services/hips2fits a service provided by CDS, the SIMBAD database, operated at CDS, Strasbourg, France (Wenger et al., 2000), NASA’s Astrophysics Data System Bibliographic Services, and the ATNF pulsar catalogue version 2.0.0 and 2.3.0. Figure 1 was made using the APLpy v2.1.0 (Robitaille & Bressert, 2012; Robitaille, 2019), which is an open-source astronomy plotting Python package.